OpenStudy (anonymous):

Please help.......How do I make a table of solutions for the equation 3x+2y=-18 using -2 and three other numbers for my x axis and using 3 and three other numbers for my y axis.Ok I was only given two numbers for x which was 2 and 0 so now I need two more numbers for x. The same with y. I was only given two numbers which was 3 and 0, so I need two numbers to complete the the table. Hope u understand

OpenStudy (anonymous):

Well, here's how I'd start: just plug in some arbitrary values for x and find the corresponding value of y, then put that into your table. So, for example, let's start with your given number x=-2. If we evaluate the expression with that assumption then we have: 3*(-2)+2y=-18 we can simplify that to read... -6+2y=-18 simplify further... 2y = -18 - (-6) 2Y = -12 Y =-6 So, the first line in your table would be X | Y -2 |6

OpenStudy (anonymous):

So how would I find the x if the y is 3?